OpenStudy (anonymous):

one year ago If you have 3.64 grams of H2, how many grams of NH3 can be produced? The equation is N2+3H2=2NH3

OpenStudy (anonymous):

41.23 g of nh3 because you do conversionsif you need the work tell me

OpenStudy (anonymous):

\[\frac{ 3.64 g H2 }{ }*\frac{ 1 mole H2 }{ 2 g H2 }*\frac{ 2 mole NH3 }{ 3 mole H2 }*\frac{ 17 g NH3 }{ 1 mole NH3 }\]